4^(-3X+1)*16^(-X)=1^(-10)
Answer by lwsshak3(11628)

4^(-3X+1)*16^(-X)=1^(-10)
**
4^(-3X+1)*16^(-X)=1^(-10)=1
take log of both sides
(-3x+1)log4+(-xlog16)=log1=0
-3xlog4+log4-xlog16=0
x(-3log4-log16)=-log4
x=-log4/-(3log4+log16)
x=.2
Check:
4^(-3X+1)*16^(-X)=1^(-10)=1
4^(-3*.2+1)*16^(.2)
4^-.4*1.7411
.5743*1.7411=1.000